. A bed mart company is in the business of manufacturing beds and pillows. The company has 40 hours for assembly and 32 hours for finishing work per day. Manufacturing of a bed requires 4 hours for assembly and 2 hours for finishing. Similarly a pillow requires 2 hours for assembly and 4 hours for finishing. Profitability analysis indicates that every bed would contribute Birr 80, while a pillow contribution is Birr 55 respectively. Find out the daily production of the company to maximise the contribution (profit). Solve the problem by graphical method.


1
Expert's answer
2022-02-28T17:49:34-0500

"x" - beds, "y" - pillows

"4x+2y\leq40\\
2x+4y\leq32\\
x\geq0\\
y\geq0\\
F(x,y)=80x+55y\mapsto max\\
4x+2y=40\\
\begin{matrix}
 x & 0&10 \\
 y & 20&0
\end{matrix}\\
2x+4y=32\\
\begin{matrix}
 x & 0&16 \\
 y & 8&0
\end{matrix}"



"x=8, y=4\\
F=80\cdot8+55\cdot4=860"